Bob did 1/3 of his homework at school. He completed 1/4 of the remaining homework at home before dinner. what fraction remained to be completed after dnner?
Answer by THANApHD(104)

so after school the remaining work will be => 1- 1/3 = 2/3
He did 1/4 of the remining work before dinner
So remaing part will be 1-1/4 = 3/4
which equals to => 2/3 x 3/4 of whole the part
= 1/2
So he must do 1/2 of his home work after dinner.